A charge, q1, of -2 micro-coulombs is located at the origin, a charge, q2, of -7 micro-coulombs is located at x = 0 cm, y = +8.3 cm, a charge, q3, of +3 micro-coulombs is located at x = 0 cm, y = -7.9 cm, and a charge, q4, of +11 micro-coulombs is located at x = 3.3 cm, y = 0 cm. What is the angle of the total electric force on q2 (charge on the +y axis) in degrees measured counter-clockwise from the +x axis?
Answer #1

F12 = kq1q2/r^2 = 9x10^9 x 2x7x10^-12 / (8.3x10^-2)^2

= 18.29 N upwards

F32 = kq1q2/r^2 = 9x10^9 x 3x7x10^-12 / ((8.3+7.9)x10^-2)^2

= 7.2 N downwards

F42 = kq1q2/r^2 = 9x10^9 x 11x7x10^-12 / (3.3^2+8.3^2)x(10^-2)^2

= 86.9N towards 4th quadrant at angle @ = tan-1(8.3/3.3) = 68.31 deg

= 86.9N @ = -68.31 deg

Adding all three vectorically in calci

F = 18.29<90 + 7.2<-90 + 86.9<-68.31

= 76.7 N @ = -65.25 deg

= 76.7N @ = 294.75 deg from +ve x axis